# DSDNet: Raw Domain Demoiréing via Dual Color-Space Synergy

### v1

@@ 3.3. Luminance Chrominance Adaptive Transformer
To effectively capture the intricate interplay between luminance and chrominance components, we propose the LCAT. Unlike conventional transformers directly processing sRGB features, LCAT dynamically integrates sRGB and YCbCr domains feature representations through an adaptive fusion mechanism. This approach enables better color fidelity by selectively emphasizing relevant chrominance information while maintaining structural consistency.
@@ 4.1. Experimental Setup
Dataset. We evaluate our method on two representative raw domain demoiréing benchmarks: the TMM22 dataset (Yue et al., 2022) and the RawVDemoiré dataset (Yue et al., 2023a). The TMM22 dataset contains 948 pairs of raw and sRGB image pairs (540 for training and 408 for testing), capturing recaptured scenes including natural images, documents, and web pages. These images are obtained using four different smartphone cameras and three display screens of varying sizes. For both training and evaluation, patches of size $256\times 256$ and $512\times 512$ are cropped to facilitate model training and fair comparison across methods. Complementing TMM22, the RawVDemoiré dataset provides a video-based benchmark, consisting of 300 raw and sRGB video pairs (250 for training and 50 for testing), each video containing 60 frames at a resolution of $1280\times 720$. Similar to TMM22, the acquisition process employs three smartphone cameras and various displays, ensuring device diversity. Following the protocol of Yue et al. (Yue et al., 2023a), we adapt this dataset for image-level demoiréing by extracting individual frames and reorganizing them into 15,000 training and 3,000 testing image pairs.
@@
Table 2. Quantitative comparison with the state-of-the-art demoiréing methods and RAW image restoration methods on RawVDemoiré dataset (Yue et al., 2023a) regarding average PSNR, SSIM, LPIPS. The best and second results are in red and blue, respectively.
@@
(Yu et al., 2022) (Dai et al., 2022) (Wang et al., 2019) (Liang et al., 2024) (Xu et al., 2024b) (Jin et al., 2023) (Yue et al., 2023a) (Song et al., 2023) (Xu et al., 2024c) (Ours)
@@
4.2. Comparison with State-of-the-Art
Quantitative comparison. We quantitatively compare the proposed method with state-of-the-art methods, including the sRGB-based methods DMCNN (Sun et al., 2018), WDNet (Liu et al., 2020b), ESDNet (Yu et al., 2022), and a raw image demoiréing method RDNet (Yue et al., 2022). We further add comparison with a raw image demoiréing method RRID (Xu et al., 2024c) with raw+sRGB input, a video demoiréing method RawVDmoiré (Yue et al., 2023a) with raw input, a raw domain low-light enhancement method DNF (Jin et al., 2023), and a dereflection method CR3Net (Song et al., 2023) with paired raw+sRGB data. For sRGB-based methods originally designed for 3-channel inputs, we implement architectural adaptations by modifying input channels to accept 4-channel Raw data and inserting bilinear upsampling layers to resolve resolution mismatches. To the RawVDemoiré dataset, we also add several sRGB-based video methods (EDVR (Wang et al., 2019), VDMoiré (Dai et al., 2022), VRT (Liang et al., 2024), and DTNet (Xu et al., 2024b)) for comprehensive comparison. To ensure a fair comparison, all baseline implementations are fine-tuned on our training dataset using their officially recommended settings. As shown in Tab. 1 and 2, our method demonstrates significant advantages on both datasets. Compared to the latest raw demoiréing methods, DSDNet achieves improvements of 0.47 dB (PSNR), 0.003 (SSIM), and 0.014 (LPIPS) over RRID (Xu et al., 2024b) on the TMM22 dataset. In addition, DSDNet even outperforms the video-based RawVDemoiré (Yue et al., 2023a) by 1.09 dB in PSNR. On the RawVDemoiré dataset, DSDNet still shows consistency superiority over the second-best method, CR3Net, by 0.78 dB.
@@ 4.3. Ablation Studies
Our proposed DSDNet integrates three key components: the YCbCr color branch, SADM, and LACT. To evaluate their individual contributions, we conduct ablation studies on the TMM22 dataset.
Effectiveness of specific modules. To evaluate the effectiveness of each module, we progressively remove the YCbCr branch, SADM, and LACT, establishing a baseline network that operates solely in the raw domain (w/o YCbCr branch). The results in Tab. 3 show that models using only raw images are limited in performance. When introducing the YCbCr branch (#2), performance consistently improves, highlighting its effectiveness in color representation. Comparing #2 (YCbCr only) with #3 (YCbCr + SADM, w/o LACT), we observe a PSNR increase from 27.63 to 27.83 dB, demonstrating that SADM enhances cross-domain interactions and improves reconstruction quality. Further incorporating LACT (#5, full model) yields a PSNR of 28.35 dB, confirming its role in luminance correction and chrominance-aware reconstruction. The full DSDNet model, integrating all three modules, achieves the highest fidelity and perceptual quality, validating the complementary benefits of the YCbCr stream, SADM, and LACT.

### v2

@@ 3.3. Luminance Chrominance Adaptive Transformer
To effectively capture the intricate interplay between luminance and chrominance components, we propose the LCAT. Unlike conventional transformers that directly process sRGB features, LCAT dynamically integrates sRGB and YCbCr domains feature representations through an adaptive fusion mechanism. This approach enables better color fidelity by selectively emphasizing relevant chrominance information while maintaining structural consistency.
@@ 4.1. Experimental Setup
Dataset. We evaluate our method on two representative raw domain demoiréing benchmarks: the TMM22 dataset (Yue et al., 2022) and the RawVDemoiré dataset (Yue et al., 2023). The TMM22 dataset contains 948 pairs of raw and sRGB image pairs (540 for training and 408 for testing), capturing recaptured scenes including natural images, documents, and web pages. These images are obtained using four different smartphone cameras and three display screens of varying sizes. For both training and evaluation, patches of size $256\times 256$ and $512\times 512$ are cropped to facilitate model training and fair comparison across methods. Complementing TMM22, the RawVDemoiré dataset provides a video-based benchmark, consisting of 300 raw and sRGB video pairs (250 for training and 50 for testing), each video containing 60 frames at a resolution of $1280\times 720$. Similar to TMM22, the acquisition process employs three smartphone cameras and various displays, ensuring device diversity. Following the protocol of Yue et al. (Yue et al., 2023), we adapt this dataset for image-level demoiréing by extracting individual frames and reorganizing them into 15,000 training and 3,000 testing image pairs.
@@
4.2. Comparison with State-of-the-Art
Quantitative comparison. We quantitatively compare the proposed method with state-of-the-art methods, including the sRGB-based methods DMCNN (Sun et al., 2018), WDNet (Liu et al., 2020b), ESDNet (Yu et al., 2022), and a raw image demoiréing method RDNet (Yue et al., 2022). We further add comparison with a raw image demoiréing method RRID (Xu et al., 2024b) with raw+sRGB input, a video demoiréing method RawVDmoiré (Yue et al., 2023) with raw input, a raw domain low-light enhancement method DNF (Jin et al., 2023), and a dereflection method CR3Net (Song et al., 2023) with paired raw+sRGB data. For sRGB-based methods originally designed for 3-channel inputs, we implement architectural adaptations by modifying input channels to accept 4-channel Raw data and inserting bilinear upsampling layers to resolve resolution mismatches. To the RawVDemoiré dataset, we also add several sRGB-based video methods (EDVR (Wang et al., 2019), VDMoiré (Dai et al., 2022), VRT (Liang et al., 2024), and DTNet (Xu et al., 2024a)) for comprehensive comparison. To ensure a fair comparison, all baseline implementations are fine-tuned on our training dataset using their officially recommended settings. As shown in Tab. 1 and 2, our method demonstrates significant advantages on both datasets. Compared to the latest raw demoiréing methods, DSDNet achieves improvements of 0.47 dB (PSNR), 0.003 (SSIM), and 0.014 (LPIPS) over RRID (Xu et al., 2024a) on the TMM22 dataset. In addition, DSDNet even outperforms the video-based RawVDemoiré (Yue et al., 2023) by 1.09 dB in PSNR. On the RawVDemoiré dataset, DSDNet still shows consistency superiority over the second-best method, CR3Net, by 0.78 dB.
Table 2. Quantitative comparison with the state-of-the-art demoiréing methods and RAW image restoration methods on RawVDemoiré dataset (Yue et al., 2023) regarding average PSNR, SSIM, LPIPS. The best and second results are in red and blue, respectively.
@@ 4.2. Comparison with State-of-the-Art
(Yu et al., 2022) (Dai et al., 2022) (Wang et al., 2019) (Liang et al., 2024) (Xu et al., 2024a) (Jin et al., 2023) (Yue et al., 2023) (Song et al., 2023) (Xu et al., 2024b) (Ours)
@@ 4.3. Ablation Studies
Our proposed DSDNet integrates three key components: the YCbCr color branch, SADM, and LCAT. To evaluate their individual contributions, we conduct ablation studies on the TMM22 dataset.
Effectiveness of specific modules. To evaluate the effectiveness of each module, we progressively remove the YCbCr branch, SADM, and LCAT, establishing a baseline network that operates solely in the raw domain (w/o YCbCr branch). The results in Tab. 3 show that models using only raw images are limited in performance. When introducing the YCbCr branch (#2), performance consistently improves, highlighting its effectiveness in color representation. Comparing #2 (YCbCr only) with #3 (YCbCr + SADM, w/o LCAT), we observe a PSNR increase from 27.63 to 27.83 dB, demonstrating that SADM enhances cross-domain interactions and improves reconstruction quality. Further incorporating LCAT (#5, full model) yields a PSNR of 28.35 dB, confirming its role in luminance correction and chrominance-aware reconstruction. The full DSDNet model, integrating all three modules, achieves the highest fidelity and perceptual quality, validating the complementary benefits of the YCbCr stream, SADM, and LCAT.
@@ 5. Conclusion
Acknowledgements.
This work is supported by the National Natural Science Foundation of China under Grants 62231018 and 62472308.